OpenStudy (anonymous):

Which number correctly completes this equation? 5 10 % × 55 = _____ Select one of the options below as your answer: A. .00275 B. .0275 C. .275 D. 2.75

OpenStudy (goformit100):

B. .0275

OpenStudy (anonymous):

5/10=0.5 0.5/100=0.005 0.005/55=0.0090909.

OpenStudy (kaiz122):

it's .005 * 55 =.275

OpenStudy (anonymous):

@DamnImBad You are correct, you just have done the wrong math at the end. 0.005 / 55 is not 0.0090909 0.005 / 55 is 0.275. Good job though! You did the right things.